In 1990,Joshua's university tuition was $1500. In 1995, tuition at the same university cost $3200. Assuming that an exponential model applies, how much will tuition cost in 2009?
Answer by checkley71(8403)

(1995-1990)/(3200-1500)=(2009-1990)/(X-1500)
5/1700=19/(X-1500
5(X-1500)=19*1700
5X-7500=32300
5X=32300+7500
5X=39800
X=39800/5
X=$7,960 IS THE PROJECTED TUITION FOR 2009.